### v2

# Detecting Misbehaviors of Large Vision-Language Models by Evidential Uncertainty Quantification

Abstract
Large vision-language models (LVLMs) have achieved substantial advances in multimodal understanding. However, when presented with challenging or distribution-shifted inputs, they frequently produce unreliable or even harmful content, such as hallucinations or toxic responses. We refer to such misalignments with human expectations as misbehaviors of LVLMs, which raise serious concerns for their deployment in critical applications. Existing research have disclosed that such misbehaviors are closely linked to model uncertainty. We find they primarily stem from two distinct sources of epistemic uncertainty: internal contradictions (conflict) and the absence of supporting information (ignorance). While existing uncertainty quantification methods typically capture only total predictive uncertainty, they struggle to distinguish between these underlying causes. To address this gap, we propose Evidential Uncertainty Quantification (EUQ), a training-free framework that explicitly decomposes epistemic uncertainty into conflict (CF) and ignorance (IG). Specifically, we interpret features from the model output head as either supporting (positive) or opposing (negative) evidence. Leveraging Dempster-Shafer Theory of belief functions, we aggregate this evidence to quantify internal conflict and knowledge gaps within a single forward pass. We extensively evaluate EUQ across four misbehavior categories, including hallucinations, jailbreaks, adversarial vulnerabilities, and out-of-distribution (OOD) failures using state-of-the-art LVLMs. Experimental results demonstrate that EUQ consistently outperforms strong baselines, achieving relative improvements of up to 10.5% in AUROC. Our evaluation further reveals that hallucinations correspond to high internal conflict and OOD failures to high ignorance. Furthermore, a layer-wise evidential uncertainty dynamics analysis provides a novel perspective on the evolution of internal representations. The source code is available at https://github.com/HT86159/EUQ.
1 Introduction
Large Vision-Language Models (LVLMs) (Liu et al., 2024c; Bai et al., 2025; Wu et al., 2024b) have demonstrated remarkable capabilities in multimodal understanding and context-aware reasoning across a variety of vision-language tasks (Ngiam et al., 2011; Chen et al., 2020). Nevertheless, their outputs can become unreliable or even harmful when faced with challenging, distribution-shifted, or adversarial inputs. Such challenges often lead to issues such as unfaithful hallucinations (Biten et al., 2022; Li et al., 2023b), security risks through jailbreaks (Qi et al., 2024; Gong et al., 2025), adversarial vulnerabilities (Fang et al., 2024; Ge et al., 2023), and failures to generalize out-of-distribution (OOD) (Yang et al., 2024; Xu et al., 2025). These misbehaviors indicate that current LVLMs are not yet fully aligned with human expectations (Bengio et al., 2025; Feng et al., 2026). As a result, such failures significantly hinder their deployment in critical applications, such as identity authentication (Li et al., 2023a), autonomous driving (Grigorescu et al., 2020), and medical diagnosis (Kumar et al., 2023). This underscores the urgent need for effective detection and mitigation methods to enhance model trustworthiness.
The connection between such misbehaviors and model uncertainty has been widely recognized (Farquhar et al., 2024; Liao et al., ; ISO, 2022). Our focus mainly lies on a significant and reducible component, epistemic uncertainty, which is the limitation in model knowledge captured by its parameters. This uncertainty has long been understood to originate from two primary sources (Denœux et al., 2020): the presence of conflicting information and the absence of supporting information. For instance, the top case in Figure 1 illustrates the former; the model correctly identifies both the text and the background image, yet their semantic inconsistency leads to a response that casts doubt on the input. In contrast, the bottom example shows the latter, with the model perceiving color and shape but expressing “cannot immediately identify” and resorting to “guessing” due to missing information.
Figure 1: The example shows how our evidential uncertainty, visualized as a token-level heatmap over the Chain-of-Thought (CoT) (Kojima et al., 2022) traces, shows that the misbehavior may stem from internal conflict and a lack of knowledge.
While misbehaviors in LVLMs often arise from internal conflicts or knowledge absence, existing uncertainty quantification (UQ) approaches focus on the total predictive uncertainty, but fail to explicitly capture such underlying causes. Most classical uncertainty quantification (UQ) methods, such as Bayesian approaches (MacKay, 1992; Blundell et al., 2015) and their variants (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017; Maddox et al., 2019), as well as internal methods of deterministic models (Sensoy et al., 2018; Malinin & Gales, 2019), are challenging to apply to LVLMs due to their substantial computational overhead. As a result, recent efforts predominantly adopt test-time sampling strategies. A typical strategy estimates overall epistemic uncertainty from token-level probabilities of a single output (Kadavath et al., 2022; Malinin & Gales, 2021), while subsequent extensions evaluate semantic variability across multiple generations (Farquhar et al., 2024; Manakul et al., 2023; Chen et al., 2025). Another line of work encourages models to verbalize their confidence (Xiong et al., ; Lin et al., ). However, such uncertainty is often unstable and uncalibrated, as LVLMs lack strong metacognitive capabilities, i.e., they struggle to reliably recognize and express their own uncertainty.
To this end, we propose Evidential Uncertainty Quantification (EUQ), which enables effective and computationally efficient detection of model misbehaviors. To the best of our knowledge, this is the first attempt to explicitly characterize two types of epistemic uncertainty in LVLMs, conflict (CF) and ignorance (IG). CF quantifies the degree of contradiction among evidence in model predictions, while IG measures the lack of information available to the model. Specifically, we draw inspiration from the interpretation of linear projection as evidence fusion (Denœux, 2019; Huang et al., 2025). Evidence is then constructed from the pre-logits features of the LVLM output head, which provide high-level signals directly linked to the model’s decisions (Zhao et al., 2024), to quantify uncertainty. We then apply basic belief assignment (BBA), which distributes belief masses over hypotheses, to convert them into evidence weights. These weights are then decomposed into positive and negative components, which represent support and contradiction to the model’s decision. The refined evidence weights are fused using Dempster’s rule of combination (Shafer, 1976), yielding CF from the conflict between positive and negative evidence and IG from the missing information in the fused evidence. As shown in Figure 1, $\mathrm{CF}$ primarily highlights concrete objects (e.g., diaper), whereas $\mathrm{IG}$ captures both objects and modifiers (e.g., vaguely). Tokens corresponding to the final decision exhibit high CF and IG values, indicating that these measures effectively capture the sources of uncertainty. We evaluate our method on misbehavior detection across four scenarios, encompassing hallucinations, jailbreaks, adversarial vulnerabilities, and OOD failures. Comprehensive experiments on DeepSeek-VL2-Tiny (Wu et al., 2024b), Qwen2.5-VL-7B (Bai et al., 2025), InternVL2.5-8B (Chen et al., 2024), and MoF-Models-7B (Tong et al., 2024) show that CF and IG consistently outperform strong baselines, achieving relative improvements of 10.4%/7.5% AUROC and 5.3%/5.5% AUPR. Empirical analysis further reveals that hallucinations correspond to high internal conflict, whereas OOD failures correspond to high ignorance. Our contributions are summarized as follows:
  • We identify that diverse misbehaviors in LVLMs primarily stem from two types of epistemic uncertainty: internal contradictions and missing supporting information. To address this, we propose a computationally efficient detection method based on Dempster-Shafer Theory (DST). It captures these fine-grained uncertainties in a single forward pass.
  • We conduct a layer-wise dynamic analysis that offers a novel perspective for interpreting the evolution of internal representations in LVLMs. This analysis also enables certain layers to distinguish among all four misbehavior categories.
  • Extensive experiments on four advanced LVLMs using our proposed Misbehavior-Bench [^1] https://huggingface.co/datasets/thuang5288/Misbehavior-Bench demonstrate that the method consistently outperforms strong baselines. It yields improvements of 10.4%/7.5% in AUROC and 5.3%/5.5% in AUPR.
2 Related Work
In this section, we first review four typical categories of misbehaviors observed in LVLMs (Section 2.1), and then discuss UQ methods that can be leveraged for detection (Section 2.2).
2.1 Misbehaviors in LVLMs
This section provides an overview of key misbehaviors observed in LVLMs, including hallucinations, jailbreaks, adversarial vulnerabilities, and failures caused by OOD inputs.
Hallucination in LVLMs denotes mismatches between visual inputs and generated text (Liu et al., 2024b). It can be categorized into three types: object hallucination, describing nonexistent objects (Biten et al., 2022; Hu et al., ; Li et al., 2023b); relation hallucination, misrepresenting spatial or semantic relations (Wu et al., 2024a); attribute hallucination, assigning wrong properties to visual entities (Liu et al., 2024a). Jailbreak refers to eliciting harmful behaviors misaligned with human intent, often triggered by visual perturbations (Carlini et al., 2023), exposing vulnerabilities beyond typical prediction errors. Such attacks are broadly categorized into optimization-based methods, which iteratively modify inputs via gradients or search strategies (Qi et al., 2024; Wang et al., 2024b; Bailey et al., 2024), and generation-based methods, which embed harmful typography on the clean images (Gong et al., 2025; Li et al., 2024a; Goh et al., 2021; Shayegani et al., ). Adversarial vulnerability in vision models stems from imperceptible adversarial perturbations that induce incorrect predictions (Szegedy et al., 2014; Li et al., 2023a; Fang et al., 2024; Ge et al., 2023; Qin et al., 2025). Recent work shows that LVLMs inherit this weakness (Sheng et al., 2021; Zhao et al., 2023; Wang et al., 2024a), remaining susceptible to visual perturbations despite their multimodal nature. OOD failure refers to the inability of a model to handle inputs outside the training distribution, challenging accurate recognition (Kim et al., 2025; Han et al., ). Prior work has focused on multimodal models, like CLIP (Radford et al., 2021), for detecting inputs outside the in-distribution (ID) (Ming et al., 2022; Jiang et al., ; Cao et al., 2024). Although OOD in LVLMs is less studied, recent work defines ID inputs as standard data and OOD inputs as style or quality shifts (Kim et al., 2025; Xu et al., 2025).
In summary, LVLMs are prone to exhibiting various misbehaviors, clearly highlighting the critical necessity of effective detection methods to ensure their reliability and robustness.
2.2 Uncertainty Quantification for LVLMs
Classical UQ methods, such as Bayesian approaches (MacKay, 1992; Blundell et al., 2015) and their variants (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017; Maddox et al., 2019), are computationally expensive and thus difficult to apply to LVLMs. Deterministic methods, such as (Malinin & Gales, 2019) and  (Sensoy et al., 2018; Li et al., 2025), the latter following an evidential framework, still require model training. In contrast, our approach performs evidence modeling and aggregation at inference, producing richer uncertainty measures without additional training, making it well suited for LVLMs. Thus, this section reviews prior work on UQ for LVLMs.
Token-wise probability-based methods estimate uncertainty within a single generation using log-likelihoods (Kadavath et al., 2022; Guerreiro et al., 2023; Duan et al., 2024) and entropy measures (Malinin & Gales, 2021). However, softmax outputs tend to be overconfident (Gal & Ghahramani, 2016; Guo et al., 2017), resulting in miscalibrated uncertainty. Sampling-based methods further estimate uncertainty by evaluating variability semantics across multiple generations. (Lin et al., 2023) estimates uncertainty via pairwise similarities and a graph Laplacian. (Farquhar et al., 2024) proposes semantic entropy to detect confabulations, utilizing external models to evaluate semantic equivalence. Other works (Raj et al., 2023; Manakul et al., 2023) design task-specific prompts and use auxiliary LLMs to assess semantic consistency. Regardless, these methods are computationally expensive due to repeated inference and heavily depend on auxiliary models. Verbal elicitation approaches, completely independent of output probabilities, estimate a model’s uncertainty by prompting it to express self-assessments in natural language. (Lin et al., ) introduces verbalization probability and demonstrates its alignment with model logits after fine-tuning. Subsequent studies (Tian et al., 2023; Zhou et al., 2023; Xiong et al., 2024) focus on prompting strategies, such as employing Chain-of-Thought (CoT) (Kojima et al., 2022) to improve verbalized uncertainty, which depends heavily on the model’s compliance with prompts (Kapoor et al., 2024).
Prior methods are often less effective at capturing the patterns of misbehaviors. In contrast, our approach leverages LVLM output head features, capturing conflict (internal contradictions) and ignorance (lack of reliable information), which enables differentiation among misbehavior types.
Figure 2: The overall framework of the proposed method applies basic belief assignment to the pre-logits feature to obtain evidence weights. These weights are then decomposed into positive and negative components, which are fused to estimate the final uncertainties that can detect different types of misbehaviors, respectively.
3 Evidential Uncertainty Quantification
This section first introduces pre-logits features in the LVLM output head and the basics of Dempster-Shafer Theory (Section 3.1). Next, these features are then interpreted as evidence for belief assignment (Section 3.2) and used to quantify conflict and ignorance via evidence fusion (Section 3.3).
3.1 Preliminary
LVLM Output Head
LVLMs typically employ an LLM with a decoder architecture, along with an output head that generally includes a projection layer and softmax for predicting the next token, as shown in Figure 2(a). The linear projection layer serves as the decision layer of LVLMs, encoding cross-modal information critical for decision making (Bi et al., 2024; Montavon et al., 2017; Zhao et al., 2024). This layer contains features directly mapped to human-readable tokens, motivating the use of the output head for uncertainty quantification. We denote the pre-logits features by $\mathbf{Z}=(z_{1},\dots,z_{I})\in\mathbb{R}^{I}$ and the output of the projection layer by $\mathbf{H}=(h_{1},\dots,h_{J})\in\mathbb{R}^{J}$, where $\mathbf{Z}$ is interpreted as evidence (Tong et al., 2021; Manchingal et al., 2025) for estimating uncertainty. Consequently, the projection layer shown in Figure 2(a) can be formalized as:
$\mathbf{H}$ $=\mathbf{Z}\mathbf{W}+\mathbf{b},$ (1)
where $\mathbf{W}\in\mathbb{R}^{I\times J}$, $\mathbf{b}\in\mathbb{R}^{I}$ denotes the weights and biases for the linear transformations, respectively.
Dempster-Shafer Theory
The Dempster–Shafer Theory (DST), also known as Evidence Theory, extends classical probability theory by providing a more flexible framework for representing and combining uncertainty derived from evidence (Dempster, 1967; Shafer, 1976) (details and illustrative examples are provided in Appendix A.2). Given a frame of discernment $\mathcal{H}$, defined as a finite set of mutually exclusive and exhaustive hypotheses, a mass function (also called a basic belief assignment, BBA) $m(\cdot)$ assigns belief to all subsets of $\mathcal{H}$. Formally, it is defined as:
$m:2^{\mathcal{H}}\rightarrow[0,1],\quad\sum_{\mathcal{S}\subseteq\mathcal{H}}m(\mathcal{S})=1;\quad m(\emptyset)=0,$ (2)
where $\mathcal{S}$ is any subset of $\mathcal{H}$, and $\emptyset$ represents the empty set. Subsets with nonzero mass are called focal sets. A mass function is simple if it assigns nonzero mass to exactly two focal sets:
$m(\mathcal{S})=s;\quad m(\mathcal{H})=1-s;\quad m(\emptyset)=0.$ (3)
DST also introduces Dempster’s rule (Shafer, 1976) for combining two mass functions $m_{1}$ and $m_{2}$, enabling multi-source evidence fusion. The rule is given by:
$\left(m_{1}\oplus m_{2}\right)(\mathcal{S})=\frac{1}{1-\kappa}\sum_{\mathcal{S}_{1}\cap\mathcal{S}_{2}=\mathcal{S}}m_{1}(\mathcal{S}_{1})m_{2}(\mathcal{S}_{2});\quad\kappa=\sum_{\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset}m_{1}(\mathcal{S}_{1})m_{2}(\mathcal{S}_{2}),$ (4)
where $(m_{1}\oplus m_{2})(\emptyset)=0$, $\mathcal{S}_{1},\mathcal{S}_{2}\subseteq\mathcal{H}$, and $\kappa$ denotes the degree of conflict between $m_{1}$ and $m_{2}$.
3.2 Belief Assignment
Due to the key role of the pre-logits feature $\mathbf{Z}$ in model predictions, we treat it as evidence for BBA. This evidence enables quantifying two primary evidential uncertainties: conflict ($\mathbf{CF}$) and ignorance ($\mathbf{IG}$). This perspective builds on the framework of (Denœux, 2019), which shows that the linear transformation can be viewed as evidence fusion of its input features via Dempster’s rule. In the remainder of this paper, we present the EUQ process based on $\mathbf{Z}$, as illustrated in Figure 2(b).
Each component $z_{i}$ of $\mathbf{Z}$ may support or contradict a candidate output feature $h_{j}$. For each pair $(z_{i},h_{j})$, we define a mass function $m_{ij}$ associated with an evidence weight $e_{ij}$, which quantifies the degree of support that $z_{i}$ provides to the validity of the feature $h_{j}$. We model the relationship between the input features and the corresponding evidence weights using an element-wise affine transformation:
$\mathbf{E}=\mathbf{A}\odot\mathbf{Z}^{\top}+\mathbf{B},$ (5)
where $\mathbf{E}\in\mathbb{R}^{I\times J}$ is the matrix of evidence weights $\{e_{ij}\}$. The parameters $\mathbf{A},\mathbf{B}\in\mathbb{R}^{I\times J}$ are obtained via closed-form estimation, as demonstrated in Lemma 1, and represent the influence of each input feature $z_{i}$ on the output feature $h_{j}$. We further decompose $\mathbf{E}$ into its positive and negative parts: $\mathbf{E}^{+}=\max(0,\mathbf{E});\;\mathbf{E}^{-}=\max(0,-\mathbf{E}),$ with entries $\{e_{ij}^{+}\}$ and $\{e_{ij}^{-}\}$, respectively. These indicate support for $h_{j}$ and its complement $\overline{\{h_{j}\}}$. Accordingly, we define positive and negative simple mass functions for each pair $(z_{i},h_{j})$ as:
$m_{ij}^{+}(\{h_{j}\})=1-\exp(-{e_{ij}^{+}}),\quad m_{ij}^{-}(\overline{\{h_{j}\}})=1-\exp(-{e_{ij}^{-}}).$ (6)
Next, we apply the Least Commitment Principle (LCP) (Smets, 1993), a conservative strategy for BBA that assigns support only to options directly justified by the available evidence. To estimate a better-calibrated weights of evidence matrix, we design the following objective under the LCP:
$\min_{\mathbf{A},\mathbf{B}}$ $\|\mathbf{A}\odot\mathbf{Z}^{\top}+\mathbf{B}\|_{2}^{2},\quad\text{s.}\text{t.}\mathbf{1}^{\top}\mathbf{B}=\mathbf{b}\cdot\mathbf{1},$ (7)
where $\mathbf{1}$ denotes the all-ones vector and $\mathbf{b}$ is the bias term of the projection layer. This constraint prevents trivial solutions and ensures equal treatment across feature dimensions.
Lemma 1 (Optimal Belief Assignment).
Given input features $\mathbf{Z}\in\mathbb{R}^{I}$ and a linear transformation with weights $W\in\mathbb{R}^{I\times J}$ and corresponding bias $b\in\mathbb{R}^{I}$, the belief assignment parameters under the Least Commitment Principle (LCP) admit the following optimal closed-form solution:
$\mathbf{A}^{*}=W-\mu_{0}(W),\quad\mathbf{B}^{*}=-\left(\mathbf{A}^{*}-\mu_{1}(\mathbf{A}^{*})\right)\odot\mathbf{Z}^{\top},$ (8)
where $\mu_{0}(\cdot)$ and $\mu_{1}(\cdot)$ compute the mean along the first and second dimensions, respectively. Here, $\mathbf{A}^{*}$ and $\mathbf{B}^{*}$ denote the optimal belief assignment parameters that minimize the commitment.
The optimality of this solution allows for a precise quantification of evidence weight, which is essential for subsequent uncertainty estimation. For full details, please refer to the AppendixA.4.
3.3 Uncertainty Estimation
We introduce the additivity of evidence weights (Lemma 2, Appendix A.5): for two simple mass functions $m_{1}(\cdot)$ and $m_{2}(\cdot)$, with associated evidence weights $e_{1}$ and $e_{2}$ respectively, if they share the same focal sets $\mathcal{S}\subseteq\mathcal{H}$, the $m_{1}\oplus m_{2}(\cdot)$ reduces to $e_{1}+e_{2}$. Formally, first-stage fusion yields:
$m(\mathcal{H})=m_{1}(\mathcal{H})\cdot m_{2}(\mathcal{H});\quad m(\mathcal{S})=1-m(\mathcal{H});\quad e=e_{1}+e_{2},$ (9)
where the $e$ is the evidence weight of $m_{1}\oplus m_{2}(\cdot)$. As a consequence, mass functions sharing the same focal sets can be directly combined, thereby alleviating the overhead of power set computation in DST (Voorbraak, 1989). This property yields the following mass functions:
$m_{j}^{+}(\{h_{j}\})=1-\exp(-{e_{j}^{+}})=1-\exp(-\sum\nolimits_{i}e_{ij}^{+});$ (10)
$m_{j}^{-}(\overline{\left\{h_{j}\right\}})=1-\exp(-{e_{j}^{-}})=1-\exp(-\sum\nolimits_{i}e_{ij}^{-}).$
Here, $\mathbf{CF}$ quantifies the conflict between the combined positive and negative evidence, while $\mathbf{IG}$ reflects the overall ignorance by aggregating all $m_{j}^{-}(\mathcal{H})$. Following the definitions of degree of conflict and ignorance in DST, these quantities are expressed as:
$\mathbf{CF}=\sum_{\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset}m^{+}(\mathcal{S}_{1})\,m^{-}(\mathcal{S}_{2}),\quad\mathbf{IG}=\sum_{j}m_{j}^{-}(\mathcal{H}),$ (11)
where $m^{+}=\bigoplus_{j}m_{j}^{+}$ and $m^{-}=\bigoplus_{j}m_{j}^{-}$ denote the combined positive and negative evidence from the second-stage fusion. Importantly, Eq. equation 11 allows computing $\mathbf{CF}$ and $\mathbf{IG}$ without enumerating the full power set of $\mathcal{H}$, avoiding the usual combinatorial explosion in DST.
Theorem 1 (Evidential Conflict and Ignorance within LVLMs).
Let $\mathbf{Z}=\{z_{i}\}_{i=1}^{I}$ denote the pre-logits feature of LVLMs, and let $m_{ij}^{k}$ be the mass function expressing the support that $z_{i}$ provides for output feature $h_{j}\in\mathcal{H}$, and $\mathcal{H}$ is the frame of discernment. The conflict $\mathbf{CF}$ and ignorance $\mathbf{IG}$ are determined by the inconsistency and insufficiency among mass functions $\{m_{ij}\}$. Specifically,
$\begin{aligned} \mathbf{CF}&=\sum\nolimits_{j}\eta^{+}_{j}\cdot\eta^{-}_{j},&\quad\mathbf{IG}&=\sum\nolimits_{j}\exp(-e_{j}^{-});\\ \eta_{j}^{+}&=\frac{\exp(e_{j}^{+})-1}{\sum_{j}\exp(e_{j}^{+})-J+1},&\quad\eta_{j}^{-}&=1-\frac{\exp(-e_{j}^{-})}{1-\prod_{j}(1-\exp(-e_{j}^{-}))}\end{aligned}$ (12)
where $\eta_{j}^{+}$ and $\eta_{j}^{-}$ denote the support and opposition ratios for component $h_{j}$, respectively. Their product measures the local conflict, and the aggregated opposition determines the overall ignorance.
Theorem 1 shows (proof in AppendixA.6), when both $\eta_{j}^{+}$ and $\eta_{j}^{-}$ are simultaneously high for the same $h_{j}$, their product becomes large, indicating a strong internal contradiction $\mathbf{CF}$. The $\mathbf{IG}$ increases as the negative evidence weights $e_{j}^{-}$ decrease, indicating higher uncertainty due to a lack of reliable information. LVLMs generate responses token by token, each with an evidential uncertainty value. We quantify the sentence-level uncertainty by averaging these values across all tokens.
4 Layer-wise Evidential Uncertainty Dynamics
This section first presents the experimental setup in Section 4.1, followed by an investigation of evidential uncertainty dynamics in LVLMs. First, we examine layer-wise dynamics to analyze how uncertainty evolves across linear layers during inference in Section 4.2. Second, we leverage the layer-wise analysis to differentiate between various misbehaviors in Section 4.3.
4.1 Experimental Settings
Table 1: Overview of datasets and evaluation types.
Scenarios Methods Size Question Type
Hallucination (Li et al., 2023b) 1000 Multiple-choice
Hallucination (Wu et al., 2024a) 1000 Multiple-choice
Jailbreak (Gong et al., 2025) 200 Open-ended
Jailbreak (Li et al., 2024b) 200 Open-ended
Jailbreak (Qi et al., 2024) 600 Open-ended
Jailbreak (Goh et al., 2021) 1800 Multiple-choice
Adversarial (Fang et al., 2024) 200 Yes-and-No
Adversarial (Ge et al., 2023) 200 Yes-and-No
OOD (Xu et al., 2025) 1300 Yes-and-No
This section summarize experimental setup. Detailed version is provided in the AppendixA.7.
Datasets We evaluate our method and baselines on hallucination scenarios using POPE (Li et al., 2023b) and R-Bench (Wu et al., 2024a), focusing on object and relation hallucinations. For jailbreak scenarios, we evaluate a range of jailbreak attacks, including FigStep (Gong et al., 2025), Hades (Li et al., 2024b), and VisualAdv (Qi et al., 2024). We further simulate typographic attacks following the protocol of (Goh et al., 2021). For adversarial scenarios, we employ two state-of-the-art attacks: ANDA (Fang et al., 2024) and PGN (Ge et al., 2023). For OOD failures, we use the dataset from (Xu et al., 2025). The aforementioned scenarios and data sources constitute our Misbehavior-Bench. All tasks include a valid, correct answer, and all provided options are reasonable.
Models We evaluate four diverse LVLMs: DeepSeek-VL2-Tiny (Wu et al., 2024b), Qwen2.5-VL-7B (Bai et al., 2025), InternVL2.5-8B (Chen et al., 2024), and MoF-Models-7B (Tong et al., 2024). These models employ varied architectures, including SwiGLU (Shazeer, 2020) and MoE (Jacobs et al., 1991). We focus on smaller models for efficiency, with scale effects analyzed in Section 5.2.
Baselines We compare against four baselines: two sampling-based methods: self-consistency (SC) (Wang et al., ), semantic entropy (SE) (Farquhar et al., 2024); and two probability-based methods—predictive entropy (PE) (Kadavath et al., 2022) and its length-normalized variant (LN-PE) (Malinin & Gales, 2021). Additionally, we include HiddenDetect (Jiang et al., 2025), originally designed for jailbreak detection, which relies on refusal cues and can also be applied to other misbehaviors.
Correctness Assessment For multiple-choice and yes/no tasks, correctness is assessed using ROUGE-L (Lin, 2004) (threshold >0.5>0.5). For open-ended tasks in jailbreak contexts, we use HarmBench’s official classifier [^2] https://github.com/centerforaisafety/HarmBench to evaluate response correctness.
Evaluation Metric for Detection We use the Area Under the ROC Curve (AUROC) to evaluate detection performance, measuring the ability to rank correct (low uncertainty) above incorrect generations. The Area Under the Precision-Recall Curve (AUPR) (Davis & Goadrich, 2006) is also reported to address data imbalance from rare misbehavior cases.
Hyper-parameters For the sampling-based methods, SC and SE, we generate exactly 10 responses per question. The temperature is set to 0.1 for the first sample. The remaining samples are drawn at 1.0 to ensure diverse generations. All experiments are conducted in NVIDIA H800 PCIe GPUs.
Figure 3: Layer-wise changes of evidential uncertainty and analysis of conflict vs. ignorance across four dataset types using Intern. Results for other models are provided in Appendix A.8.
Figure 4: Density distributions of $\mathbf{CF}$, $\mathbf{IG}$, and entropy for each type of misbehavior in Intern, comparing the target misbehavior against others. Results for other models are provided in Appendix A.8.
4.2 Layer-wise Dynamics of Conflict and Ignorance
EUQ enables uncertainty quantification at every linear layer of decoder blocks, allowing us to investigate the evolving trends of $\mathbf{CF}$ and $\mathbf{IG}$ across the entire decoder.
Observation 1.
Across decoder layers, ignorance tends to decrease while conflict increases.
[…]
4.3 Distinguishing Misbehaviors via Evidential uncertainty
From Figure 3, certain decoder layers exhibit clear distinctions in their uncertainty curves across misbehaviors. Motivated by this, we leverage layer-wise $\mathbf{CF}$ and $\mathbf{IG}$ to distinguish different misbehaviors. We conducted one-vs-rest density comparisons, where each misbehavior is contrasted with the others under these three metrics. Figure 4 presents the resulting distributions of $\mathbf{CF}$, $\mathbf{IG}$, and entropy across the four misbehavior types. A clear pattern emerges: $\mathbf{CF}$ and $\mathbf{IG}$ produce clearer separations between distributions than PE, highlighting their discriminative ability. Among the four types, adversarial examples are the most distinguishable, as their density curves deviate sharply from the others due to the pronounced distributional shift caused by pixel-level perturbations. These results provide further empirical evidence that epistemic uncertainty, arising from conflict and information gaps, though currently the separation is apparent only in certain decoder layers.
5 Detection Performance Analysis
Table 2: Accuracy results of DeepSeek, Qwen, Intern, and MoF across the four misbehavior types. Best and next-best results are marked in bold and underlined, respectively.
[…]
Hallucination 0.710 0.732 0.688 0.456 0.647
Jailbreak 0.895 0.893 0.718 0.567 0.768
Adversarial 0.208 0.503 0.607 0.236 0.389
OOD 0.346 0.272 0.456 0.731 0.451
Table 3: Average AUROC and AUPR of all methods across LVLMs and datasets.
Method AUROC AUPR
SC 0.626 0.730
SE 0.624 0.661
PE 0.701 0.656
LN-PE 0.704 0.660
HiddenDetect 0.707 0.658
CF (ours) 0.812 0.783
IG (ours) 0.783 0.785
Table 4: The AUROC and AUPR of our methods and baselines on DeepSeek-VL2 (DeepSeek), Qwen2.5-VL (Qwen), InternVL2.5 (Intern), and MoF-Models (MoF) in adversarial, OOD, hallucination, and jailbreak settings. Best and next-best results are marked in bold and underlined, respectively.
[…]
AUROC AUPR AUROC AUPR AUROC AUPR AUROC AUPR AUROC AUPR
Hallucination datas from (Li et al., 2023b) and (Wu et al., 2024a).
SC 0.660 0.734 0.640 0.815 0.696 0.883 0.500 0.758 0.624 0.798
SE 0.517 0.649 0.501 0.554 0.775 0.582 0.722 0.510 0.629 0.574
PE 0.771 0.574 0.742 0.741 0.755 0.634 0.701 0.619 0.742 0.642
LN-PE 0.758 0.570 0.574 0.576 0.755 0.634 0.702 0.619 0.697 0.600
HiddenDetect 0.594 0.528 0.792 0.570 0.590 0.523 0.827 0.845 0.703 0.614
CF(ours) 0.774 0.781 0.802 0.835 0.611 0.843 0.855 0.838 0.761 0.824
IG(ours) 0.716 0.533 0.591 0.745 0.768 0.898 0.553 0.757 0.657 0.733
Jailbreak attacks from (Gong et al., 2025), (Li et al., 2024b), (Qi et al., 2024), and (Goh et al., 2021).
SC 0.606 0.606 0.512 0.861 0.546 0.781 0.920 0.846 0.646 0.774
SE 0.643 0.746 0.537 0.790 0.623 0.881 0.869 0.532 0.668 0.737
PE 0.564 0.633 0.757 0.890 0.716 0.731 0.852 0.503 0.722 0.689
LN-PE 0.657 0.561 0.703 0.891 0.725 0.698 0.853 0.893 0.735 0.761
HiddenDetect 0.842 0.746 0.842 0.802 0.623 0.543 0.699 0.586 0.752 0.669
CF(ours) 0.844 0.791 0.535 0.748 0.762 0.739 0.886 0.534 0.757 0.703
IG(ours) 0.673 0.795 0.541 0.749 0.585 0.711 0.859 0.860 0.665 0.779
Adversarial examples from (Fang et al., 2024) and (Ge et al., 2023).
SC 0.739 0.633 0.660 0.746 0.606 0.729 0.593 0.778 0.650 0.722
SE 0.669 0.838 0.688 0.514 0.634 0.557 0.552 0.707 0.636 0.654
PE 0.621 0.604 0.701 0.518 0.701 0.524 0.674 0.587 0.674 0.558
LN-PE 0.792 0.574 0.702 0.518 0.700 0.524 0.674 0.587 0.717 0.551
HiddenDetect 0.646 0.532 0.802 0.737 0.672 0.637 0.591 0.887 0.678 0.698
CF(ours) 0.921 0.928 0.847 0.738 0.706 0.773 0.868 0.832 0.836 0.818
IG(ours) 0.976 0.787 0.767 0.713 0.702 0.774 0.999 0.856 0.861 0.783
OOD inputs from (Xu et al., 2025).
SC 0.557 0.567 0.663 0.650 0.528 0.514 0.590 0.774 0.585 0.626
SE 0.526 0.622 0.592 0.698 0.622 0.659 0.505 0.736 0.561 0.679
PE 0.690 0.794 0.779 0.896 0.564 0.623 0.630 0.620 0.666 0.733
LN-PE 0.689 0.793 0.786 0.885 0.563 0.612 0.630 0.620 0.667 0.728
HiddenDetect 0.677 0.670 0.776 0.534 0.729 0.621 0.594 0.778 0.694 0.651
CF(ours) 0.809 0.572 0.996 0.994 0.791 0.651 0.979 0.930 0.894 0.787
IG(ours) 0.999 0.963 0.997 0.701 0.795 0.866 0.999 0.855 0.948 0.846
5.1 Misbehaviors Detection
Before evaluating misbehavior detection, we report the accuracy of four LVLMs across the prepared datasets. As shown in Table 3, adversarial examples yield the lowest accuracy, followed by OOD inputs. Jailbreak samples show the highest accuracy except on MMBench, likely due to LVLMs recognizing and refusing most jailbreak prompts. We begin by evaluating our method and baseline approaches across four distinct data types that elicit varied misbehaviors: hallucinated data, jailbreak attacks, adversarial examples, and OOD inputs. As shown in Table 3, CF/IG outperforms the best baseline by 10.4%/7.5% AUROC and 5.3%/5.5% AUPR on average across all models and misbehavior types. As shown in Table 12, $\mathbf{CF}$ consistently achieves superior detection performance for hallucinations, with average AUROC and AUPR scores of 0.761 and 0.824, respectively, outperforming all baselines.
Observation 2.
Hallucinations are more easily detected by conflict ($\mathbf{CF}$), whereas OOD failures are more effectively captured by ignorance ($\mathbf{IG}$).
Moreover, for jailbreak and adversarial examples, $\mathbf{CF}$ and $\mathbf{IG}$ achieve comparable results. These observations suggest that hallucinations are more likely caused by internal conflicts within the model, whereas OOD failures primarily arise from a lack of relevant information. Importantly, while HiddenDetect was designed for jailbreak detection, CF marginally outperforms it on that task and exhibits a larger advantage on other detection tasks. Specifically, CF improves over HiddenDetect by 0.5% in AUROC and 0.4% in AUPR. Our $\mathbf{CF}$, $\mathbf{IG}$, and entropy-based methods outperform sampling-based approaches, suggesting that relying solely on output consistency is insufficient to capture the model’s internal cognitive issues in misbehaviors. Moreover, our approach attains competitive performance with substantially lower computational overhead, highlighting its efficiency and practicality.
5.2 Ablation Study
Figure 5: Ablation study on temperature (left) and model scale (right) across all datasets using Intern.
Table 5: None-of-the-above rates (%) of two LVLMs in the hallucination scenario.
Model NoA (option) NoA (prompted)
Qwen 0.27 4.93
Intern 0.00 0.53
We perform ablation studies to examine the effect of model scale and feature layers on our method. We also conduct an efficiency analysis to measure computation and inference latency.
Temperature We examine the effect of temperature on LVLM generation (Figure 5, left), evaluating eight settings from $0.1$ to $1.4$. Both $\mathbf{CF}$ and $\mathbf{IG}$ remain stable, suggesting robustness of our method to this hyperparameter.
Model Size The right panel of Figure 5 compares models with 4B, 8B, 26B, and 38B parameters to illustrate the effect of scale. Detection performance is strong for the 4B and 38B models. Small models produce obvious errors that are easily captured, medium models generate subtler, less detectable errors, and large models produce mostly correct outputs, making the remaining misbehaviors easier to detect. Additionally, to verify that our method can still capture Observations 1 and 2 on a larger-scale model (e.g., 72B), we conducted experiments and report the results in Appendix A.8.
Prompting for Abstention We tested external prompting by adding a "None of the above" option (Wang & Nalisnick, 2025). As shown in Table 5, the models rarely selected it. Qwen chose it only 0.27% of the time, and Intern 0.00%. Even after reinforcing the instruction with If you are unsure, please select "None of the above"., the selection rates remained extremely low, with Qwen choosing it 4.39% of the time and Intern 0.53%, indicating persistent overconfidence in multiple-choice scenarios. This shows that prompt-based strategies have limited ability to elicit uncertainty, underscoring the need for methods that do not rely on prompting.
[…]
Table 6: Comparison of AUROC and average runtime per example across Intern.
Method Model Inference SC SE PE LN-PE HiddenDetect CF IG
Time (s) $9.6{\times}10^{-2}$ $8.9{\times}10^{-1}$ $9.0{\times}10^{-1}$ $3.1{\times}10^{-6}$ $6.1{\times}10^{-6}$ $2.0{\times}10^{-2}$ $9.1{\times}10^{-4}$ $4.5{\times}10^{-3}$
AUROC 0.626 0.624 0.701 0.704 0.707 0.812 0.783
6 Discussion
Relation with EDL-based methods
Existing applications of evidence theory in LVLMs (e.g., (Li et al., 2025; Ma et al., 2025)) are typically built on evidential deep learning (EDL) (Sensoy et al., 2018), which follows a paradigm fundamentally different from ours. These approaches are grounded in Subjective Logic (SL) and require explicit training or fine-tuning, which limits scalability to large-scale models. In contrast, EUQ focuses on detecting misbehavior without additional training and directly leverages the full expressive form of DST rather than the SL formulation. We hope this work broadens the perspective on DST-based methods and highlights an alternative evidential direction for deep learning models and LVLMs. Further details are given in Appendix A.8.5
Scope and Applicability
Our method interprets linear transformations as evidence fusion operators, which allows EUQ to apply to any model with a linear projector, including architectures such as BERT, ResNet, and LLMs. This generality extends beyond VLMs, and Appendix A.8.3 provides a toy example on convolutional networks to illustrate this. While requiring access to internal representations limits its use with closed-source APIs like GPT-4 (Achiam et al., 2023), it provides valuable fine-grained signals for failure diagnosis and model improvement.
7 Conclusion
In this work, we categorize the typical misbehaviors of LVLMs, including hallucinations, jailbreaks, adversarial vulnerabilities, and OOD failures. To detect and distinguish these misbehaviors, we introduce Evidential Uncertainty Quantification (EUQ), the first attempt to explicitly characterize two types of epistemic uncertainty in LVLMs. Furthermore, EUQ can be leveraged to interpret the internal evolution of the model decoder: ignorance generally decreases while conflict increases. Additionally, hallucination cases are primarily characterized by high internal conflict, whereas OOD failures mainly result from a lack of information. Experiments on four LVLMs show that EUQ consistently improves AUROC and AUPR, suggesting evidential reasoning as a promising direction for fine-grained uncertainty quantification, model interpretation, and misbehavior identification.
Acknowledgments
This work was supported by the National Key Research and Development Program (Grant No. 2024YFE0202900), the National Natural Science Foundation of China (Grant Nos. 62436001, 62536001, and 62406021), and the Joint Foundation of Ministry of Education for Innovation team (Grant No. 8091B042235).
Ethics Statement
This work studies misbehavior detection in LVLMs, including behaviors that may generate harmful content. Our experiments are controlled and do not involve real users. The goal is to improve model safety and reliability, mitigating potential harm from such behaviors.
Reproducibility Statement
All methods, models (with version numbers), datasets, and experimental settings are fully described to ensure reproducibility. This includes the implementation of our approach, hyperparameters, evaluation metrics, and baseline comparisons.
Appendix A Appendix
A.1 Overview
The Appendix provides supplementary material to support and extend the main content of the paper. We begin in Subsection A.2 with the theoretical foundation of Dempster–Shafer Theory, which forms the basis of our approach. Subsections A.4 to A.6 present complete proofs for Lemma 1, Lemma 2, and Theorem 1, respectively. Subsection A.7 details the experimental configurations, while Subsection A.8 offers additional results that further validate our method. Finally, Subsection A.9 outlines our usage of large language models in this work.
A.2 Dempster-Shafer Theory Foundation
The Dempster-Shafer Theory (DST), proposed by Dempster (Dempster, 1967) and Shafer (Shafer, 1976), generalizes classical probability theory to manage uncertainty and partial belief. It employs basic belief assignments (BBA) that distribute belief among subsets of the frame of discernment. This allows for a fine-grained representation of uncertainty compared to the traditional probability, which assigns specific probabilities to individual events (i.e., elements within the frame). This theory can fuse evidence from different sources using Dempster’s rule of combination. Below, we recall the key definitions employed throughout the main paper.
Mass Function
Let $\mathcal{H}=\{h_{1},h_{2},\dots,h_{J}\}$ represent the frame of discernment, which contains $J$ possible outcomes. In this context, a mass function $m(\cdot)$ maps subsets of the frame of discernment $2^{\mathcal{H}}$ to the interval $[0,1]$, indicating the degree of belief assigned to each subset. The mass function is subject to the normalization condition,
$\sum_{\mathcal{S}\subseteq\mathcal{H}}m(\mathcal{S})=1;\quad m(\emptyset)=0,$ (13)
where $\mathcal{S}$ is any subset of $\mathcal{H}$, and $\emptyset$ represents the empty set.
Focal Set
For a subset $\mathcal{S}\subseteq\mathcal{H}$, if $m(\mathcal{S})>0$, $\mathcal{S}$ is called a focal set of $m(\cdot)$.
Simple Mass Function
Specifically, a mass function is called simple when it assigns belief exclusively to one specific subset $\mathcal{S}\subseteq\mathcal{H}$ and the $\mathcal{H}$. Formally, it is defined as follows:
$m(\mathcal{S})=s;\quad m(\mathcal{H})=1-s,$ (14)
where $\mathcal{S}\neq\emptyset$ and $s\in[0,1]$ represents the degree of support for A. In particular, the mass $m(\mathcal{H})$, assigned to the entire frame, commonly indicates the degree of ignorance, as it exhibits no preferential allocation towards any particular subset.
Dempster’s Rule of Combination
Given two mass functions, $m_{1}(\cdot)$ and $m_{2}(\cdot)$, which represent evidence from two different sources (e.g., agents), the combined mass function for all $\mathcal{S}\subseteq\mathcal{H}$, with $\mathcal{S}\neq\emptyset$, is computed through Dempster’s rule of combination (Shafer, 1976) as follows:
$\left(m_{1}\oplus m_{2}\right)(\mathcal{S})=\frac{1}{1-\kappa}\sum_{\mathcal{S}_{1}\cap\mathcal{S}_{2}=\mathcal{S}}m_{1}(\mathcal{S}_{1})m_{2}(\mathcal{S}_{2});\quad\kappa=\sum_{\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset}m_{1}(\mathcal{S}_{1})m_{2}(\mathcal{S}_{2}),$ (15)
where $\left(m_{1}\oplus m_{2}\right)(\emptyset)=0$, and the $\kappa$ serves as an important metric to measure the degree of conflict between $m_{1}(\cdot)$ and $m_{2}(\cdot)$.
Belief and Plausibility Functions
Given a mass $m(\cdot)$, two useful functions, the belief and plausibility functions, are defined, respectively, as
$Bel(\mathcal{S}_{1})=\sum_{\mathcal{S}_{2}\subseteq\mathcal{S}_{1}}m(\mathcal{S}_{2});\quad Pl(\mathcal{S}_{1})=\sum_{\mathcal{S}_{2}\cap\mathcal{S}_{1}\not=\emptyset}m(\mathcal{S}_{2}).$ (16)
The belief function $Bel(\mathcal{S}_{1})$ represents the degree of certainty that the true state lies within the subset $\mathcal{S}_{1}$ based on all available evidence, excluding any possibility outside of $\mathcal{S}_{1}$. In contrast, the plausibility function $Pl(\mathcal{S}_{1})$ indicates the degree of belief that the true state may lie within $\mathcal{S}_{1}$, without ruling out possibilities.
Contour Function
In the case of singletons (only one element in a subset, e.g., $\{h_{1}\}$), the plausibility function $Pl(\cdot)$ is restricted to the contour function $pl(\cdot)$ (i.e., $pl(h_{j})=Pl(\{h_{j}\}),\;\forall h_{j}\in\mathcal{H}$). The contour function $pl(h_{j})$ measures the plausibility of each singleton hypothesis and assesses the uncertainty of each possible outcome independently. Furthermore, given two contour functions $pl_{1}(\cdot)$ and $pl_{2}(\cdot)$, associated with mass functions $m_{1}(\cdot)$ and $m_{2}(\cdot)$ respectively, they can be combined as
$pl_{1}\oplus pl_{2}(h_{j})=\frac{1}{1-\kappa}pl_{1}(h_{j})pl_{2}(h_{j}),\quad\forall h_{j}\in\mathcal{H}.$ (17)
This combination rule simplifies evidence aggregation by directly multiplying the plausibilities of singletons, making the process more efficient.
A.3 Illustrative Example
To build intuition for DST, we present a simple illustrative example. Consider a light-bulb switch whose state is either On or Off, giving the hypothesis space $\mathcal{H}=\{\text{On},\,\text{Off}\}$. In classical probability, probabilities are assigned only to the individual states:
$P(\text{On})+P(\text{Off})=1.$
If $P(\text{On})=P(\text{Off})=0.5$, the model cannot tell whether this means that the bulb is truly equally likely to be On or Off, or simply that we do not know its state. DST addresses this limitation by introducing a Basic Belief Assignment (BBA): $m(\cdot)$ defined over the power set $2^{\mathcal{H}}$:
$\sum_{S\subseteq\mathcal{H}}m(S)=1,\quad m(\emptyset)=0.$
Here, $S$ may be a single state (e.g., $\{\text{On}\}$) or the full set $\mathcal{H}$$\textbf{On},\textbf{Off}$. This allows for $2^{|\mathcal{H}|}-1=3$ distinct mass assignments, enabling richer uncertainty representation. Importantly, the belief assigned to the full set directly quantifies ignorance, i.e., how much uncertainty we have about whether the bulb is On or Off.
Conflict and ignorance from three observers
Consider again the frame $\mathcal{H}=\{\text{On},\text{Off}\}$. Three independent observers provide BBAs describing the state of the light bulb:
$m_{1}(\{\text{On}\})$ $=0.3,\quad m_{1}(\{\text{Off}\})=0.2,\quad m_{1}(\mathcal{H})=0.5,$
$m_{2}(\{\text{On}\})$ $=0.6,\quad m_{2}(\{\text{Off}\})=0.2,\quad m_{2}(\mathcal{H})=0.2,$
$m_{3}(\{\text{On}\})$ $=0.1,\quad m_{3}(\{\text{Off}\})=0.8,\quad m_{3}(\mathcal{H})=0.1.$
The belief each observer assigns to the full set $\mathcal{H}$ quantifies ignorance:
$IG_{1}$ $=m_{1}(\mathcal{H})=0.5,\quad IG_{2}=m_{2}(\mathcal{H})=0.2,\quad IG_{3}=m_{3}(\mathcal{H})=0.1.$
To combine two sources of evidence, DST uses Dempster’s rule of combination, formally expressed as:
$K=\sum_{B\cap C=\emptyset}m_{1}(B)\,m_{2}(C),$ (18)
where the sum is over all pairs of mutually exclusive subsets $B$ and $C$. Here, $K$ quantifies the conflict between the two BBAs.
For this example:
$K_{12}$ $=m_{1}(\{\text{On}\})\,m_{2}(\{\text{Off}\})+m_{1}(\{\text{Off}\})\,m_{2}(\{\text{On}\})$
$=0.3\cdot 0.2+0.2\cdot 0.6=0.18,$
$K_{23}$ $=m_{2}(\{\text{On}\})\,m_{3}(\{\text{Off}\})+m_{2}(\{\text{Off}\})\,m_{3}(\{\text{On}\})$
$=0.6\cdot 0.8+0.2\cdot 0.1=0.5.$
As shown, $K_{23}>K_{12}$, indicating that $m_{2}$ and $m_{3}$ exhibit stronger disagreement than $m_{1}$ and $m_{2}$, resulting in a higher conflict value.
A.4 Proof of Lemma 1
Preliminary
LVLMs typically use an LLM with a decoder architecture to predict the next token conditioned on vision-language features. To avoid overconfidence (Jiang et al., 2024) and achieve more precise uncertainty quantification, we focus on pre-logits features from the LVLM. These features mainly represent rich vision-language perceptual information (Basu et al., ; Bi et al., 2024) and play a key role in decision making of LVLMs (Montavon et al., 2017; Zhao et al., 2024). Specifically, in an linear projector layer, We denote the pre-logits features by $\mathbf{Z}=(z_{1},\dots,z_{I})\in\mathbb{R}^{I}$ and the output of the projection layer by $\mathbf{H}=(h_{1},\dots,h_{J})\in\mathbb{R}^{J}$, where $\mathbf{Z}$ is interpreted as evidence (Tong et al., 2021; Manchingal et al., 2025) for estimating uncertainty. Consequently, the projection layer shown in Figure 2(a) can be formalized as:
$\mathbf{H}$ $=\mathbf{Z}\mathbf{W}+\mathbf{b},$ (19)
where $\mathbf{W}\in\mathbb{R}^{I\times J}$, $\mathbf{b}\in\mathbb{R}^{I}$ denotes the weights and biases for the linear transformations, respectively.
Due to the key role of the pre-logits feature $\mathbf{Z}$ in model decisions, we treat it as evidence for belief assignment. This evidence enables quantifying two primary evidential uncertainties: conflict ($\mathbf{CF}$) and ignorance ($\mathbf{IG}$). This perspective is grounded in the theoretical framework of (Denœux, 2019), which demonstrates that the output of an FFN can be interpreted as the combination of simple mass functions derived from its input features via Dempster’s rule of combination. In the remainder of this paper, we detail the EUQ process based on the FFN feature $\mathbf{Z}$.
Each component $z_{i}$ of $\mathbf{Z}$ may support or contradict a candidate output feature $h_{j}$. For each pair $(z_{i},h_{j})$, we define a mass function $m_{ij}$ associated with an evidence weight $e_{ij}$, which quantifies the degree of support that $z_{i}$ provides to the validity of the feature $h_{j}$. We model the relationship between the input features and the corresponding evidence weights using an affine transformation:
$\mathbf{E}=\mathbf{A}\odot\mathbf{Z}^{\top}+\mathbf{B},$ (20)
where $\mathbf{E}\in\mathbb{R}^{I\times J}$ is the matrix of evidence weights, and $\mathbf{A},\mathbf{B}\in\mathbb{R}^{I\times J}$ are parameter matrices.
To ensure that belief is only assigned when sufficiently supported by evidence, we adopt the Least Commitment Principle (LCP) (Smets, 1993), which minimizes unwarranted assumptions. Under this principle, the optimal evidence weights are obtained by solving the following regularized optimization problem:
$\min_{\mathbf{A},\mathbf{B}}\quad\left\|\mathbf{A}\odot\mathbf{Z}^{\top}+\mathbf{B}\right\|_{2}^{2},\quad\text{subject to}\quad\mathbf{1}^{\top}\mathbf{B}=b\cdot\mathbf{1},$ (21)
where $\mathbf{1}$ denotes the all-ones vector, and $b$ is the bias term in the linear transformation that regulates the global evidence level across hypotheses. This constraint enforces cautious belief assignment under the LCP.
Proof Outline.
We outline the main steps as follows:
Step 1: Reformulate the optimization objective in terms of scalar parameters $\alpha_{ij}$, $\beta_{ij}$, and center the input $z_{ni}$ to simplify the expressions.
Step 2: Rewrite the loss function using centered variables to eliminate cross terms and reduce it to a sum of squares in $\alpha_{ij}$ and shifted $\beta_{ij}^{\prime}$.
Step 3: Solve for the optimal $\alpha_{ij}^{*}$ under a constraint that ensures the sum of $\beta_{ij}$ matches the bias term $\mathbf{b}_{j}$.
Step 4: Recover $\beta_{ij}^{*}$ by adjusting for centering and express the final solution in closed matrix form for $\mathbf{A}^{*}$ and $\mathbf{B}^{*}$. ∎
Proof.
We begin by rewriting the original problem in its component-wise form:
$\min_{{\alpha}_{ij},\beta_{ij}}$ $\sum_{n,i,j}\left(\alpha_{ij}\cdot{z}_{ni}+\beta_{ij}\right)^{2},\quad\text{s.}\text{t.}\;\sum_{i}\beta_{ij}=\mathbf{b}_{j},$ (22)
where $\{\alpha_{ij}\}$ and $\{\beta_{ij}\}$ denote the individual components of the matrices $\mathbf{A}\in\mathbb{R}^{I\times J}$ and $\mathbf{B}\in\mathbb{R}^{I\times J}$, respectively, while $\{z_{ni}\}$ and $\{\mathbf{b}_{j}\}$ are the components of the vectors $\mathbf{Z}\in\mathbb{R}^{I}$ and $\mathbf{b}\in\mathbb{R}^{J}$, respectively. Here, $n$ denotes the number of tokens generated in a single output sequence. For convenience in the subsequent analysis, we first center the variable $z_{ni}$ by defining
$z^{\prime}_{ni}=z_{ni}-\mu_{i},$ (23)
where $\mu_{i}=\frac{1}{N}\sum_{n}z_{ni}$ and $z^{\prime}_{ni}$ denotes the centered version of $z_{ni}$, defined as $z^{\prime}_{ni}=z_{ni}-\mu_{i}$. By substituting $z_{ni}$ with $z^{\prime}_{ni}+\mu_{i}$, the objective function becomes:
$\sum_{n,i,j}\left(\alpha_{ij}\cdot{z}^{\prime}_{ni}+\beta_{ij}+\alpha_{ij}\cdot\mu_{i}\right)^{2}=\sum_{n,i,j}\left(\alpha_{ij}\cdot{z}^{\prime}_{ni}+\beta^{\prime}_{ij}\right)^{2}$ (24)
$=$ $\sum_{n,i,j}\alpha_{ij}^{2}{z}^{\prime 2}_{ni}+2\alpha_{ij}\beta_{ij}{z}^{\prime}_{ni}+\beta_{ij}^{\prime 2}=\sum_{n,i,j}\alpha_{ij}^{2}{z}^{\prime 2}_{ni}+\beta_{ij}^{\prime 2}+\sum_{i,j}2\alpha_{ij}\beta_{ij}\underbrace{\sum_{n}{z}^{\prime}_{ni}}_{0}$
$=$ $\sum_{n,i,j}\alpha_{ij}^{2}{z}^{\prime 2}_{ni}+\beta_{ij}^{\prime 2},$
where $\beta_{ij}^{\prime}=\beta_{ij}+\alpha_{ij}\cdot\mu_{i}$. Next, we proceed to compute the optimal estimate of $\alpha_{ij}$, denoted as $\alpha_{ij}^{*}$. Furthermore, the objective function can be expressed as
$\sum_{n,i,j}\alpha_{ij}^{2}{z}^{\prime 2}_{ni}+\beta_{ij}^{\prime 2}=\sum_{i}\left(\sum_{n}{z}^{\prime 2}_{ni}\right)\left(\sum_{j}\alpha_{ij}^{2}\right)+\beta_{ij}^{\prime 2}$ (25)
Consequently, it satisfies the following constraint:
$\sum_{i}\beta^{\prime}_{ij}=\mathbf{b}_{j}^{\prime}=\mathbf{b}_{j}+\sum_{i}\alpha_{ij}\cdot\mu_{i},$ (26)
This leads to the estimate
$\alpha_{ij}^{*}=\hat{\alpha}_{ij}-\frac{1}{J}\sum_{j}\hat{\alpha}_{ij};\quad\beta_{ij}^{\prime*}=\frac{1}{J}\mathbf{b}_{j}^{\prime}=\frac{1}{J}(\mathbf{B}{j}+\sum_{i}\alpha^{*}_{ij}\cdot\mu_{i}),$ (27)
where $\hat{\alpha}_{ij}$ denotes the maximum likelihood estimate of $\alpha_{ij}$, corresponding to the model parameter $w_{ij}$ in $\mathbf{W}$. We then derive a closed-form expression for $\beta_{ij}^{*}$ as follows:
$\beta_{ij}^{*}$ $=\beta_{ij}^{\prime*}-\alpha_{ij}^{*}\cdot\mu_{i}=\frac{1}{J}\left(\mathbf{b}_{j}+\sum_{i}\alpha_{ij}^{*}\cdot\mu_{i}\right)-\alpha_{ij}^{*}\cdot\mu_{i}$ (28)
$=\frac{1}{J}\mathbf{b}_{j}-\left(\frac{1}{J}\sum_{i}\alpha_{ij}^{*}-\alpha_{ij}^{*}\right)\cdot\mu_{i}.$
Since most components of $\mathbf{b}_{j}$ in LVLMs are close to zero, we omit this term for simplicity. The final expressions for the optimal estimates are then given by:
$\alpha_{ij}^{*}=\hat{\alpha}_{ij}-\frac{1}{J}\sum_{j}\hat{\alpha}_{ij},\quad\beta_{ij}^{*}=-\left(\frac{1}{J}\sum_{i}\alpha_{ij}^{*}-\alpha_{ij}^{*}\right)\cdot\mu_{i}.$ (29)
These expressions can be compactly written in matrix form as:
$\mathbf{A}^{*}=W-\mu_{0}(W),\quad\mathbf{B}^{*}=-\left(\mathbf{A}^{*}-\mu_{1}(\mathbf{A}^{*})\right)\odot\mathbf{Z}^{\top},$ (30)
where $\mu_{0}(\cdot)$ and $\mu_{1}(\cdot)$ denote the mean over columns and rows, respectively, and $\odot$ denotes the element-wise product. ∎
A.5 Proof of Lemma 2
Lemma 2 (Additivity of Evidence Weights (Dempster, 1967)).
Let $m_{1}$ and $m_{2}$ be two simple mass functions defined over the same focal set $\mathcal{S}\subseteq\mathcal{H}$, with associated evidence weights $e_{1}$ and $e_{2}$, respectively. Under Dempster’s rule of combination, the resulting mass function $m=m_{1}\oplus m_{2}$ remains simple and retains $\mathcal{S}$ as its focal set. The corresponding weight of evidence is subsequently defined as:
$m(\mathcal{H})=m_{1}(\mathcal{H})\cdot m_{2}(\mathcal{H});\quad m(\mathcal{S})=1-m(\mathcal{H});\quad e=e_{1}+e_{2}.$ (31)
Proof Outline.
We outline the key steps as follows:
Step 1: Represent the simple mass functions $m_{1}$ and $m_{2}$ over the same focal set $\mathcal{S}$, and express their evidence weights $e_{1}$, $e_{2}$.
Step 2: Apply Dempster’s rule of combination with zero conflict $\kappa=0$, to obtain the combined mass function.
Step 3: Express the combined evidence weight $e$ in terms of $e_{1}$ and $e_{2}$, showing additivity $e=e_{1}+e_{2}$. ∎
Proof.
Since both $m_{1}(\cdot)$ and $m_{2}(\cdot)$ are simple mass functions that share the same focal set $\mathcal{S}$:
$m_{1}(\mathcal{S})=s_{1},$ $m_{1}(\mathcal{H})=1-s_{1};\quad e_{1}=-\ln(1-s_{1})$ (32)
$m_{2}(\mathcal{S})=s_{2},$ $m_{2}(\mathcal{H})=1-s_{2};\quad e_{2}=-\ln(1-s_{2}).$
Applying Dempster’s rule of combination equation 15, we compute the combined mass function as:
$(m_{1}\oplus m_{2})(\mathcal{S})$ $=\frac{1}{1-\kappa}\left[m_{1}(\mathcal{S})\cdot(m_{2}(\mathcal{S})+m_{2}(\mathcal{H}))+m_{1}(\mathcal{H})\cdot m_{2}(\mathcal{S})\right]$ (33)
$=\frac{1}{1-\kappa}\left[s_{1}+(1-s_{1})\cdot s_{2}\right],$
$(m_{1}\oplus m_{2})(\mathcal{H})$ $=\frac{1}{1-\kappa}\cdot m_{1}(\mathcal{H})\cdot m_{2}(\mathcal{H})$
$=\frac{1}{1-\kappa}\cdot(1-s_{1})(1-s_{2}),$
where the conflict mass $\kappa=\sum_{\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset}m_{1}(\mathcal{S}_{1})m_{2}(\mathcal{S}_{2})=0$, since both mass functions share the same focal set $\mathcal{S}$. Therefore, the expressions simplify to:
$(m_{1}\oplus m_{2})(\mathcal{S})$ $=s_{1}+(1-s_{1})\cdot s_{2}=1-(1-s_{1})(1-s_{2}),$ (34)
$(m_{1}\oplus m_{2})(\mathcal{H})$ $=(1-s_{1})(1-s_{2}).$
Accordingly, the evidence weight of the combined mass function $(m_{1}\oplus m_{2})(\cdot)$ is given by:
$e$ $=-\ln\left((1-s_{1})(1-s_{2})\right)$ (35)
$=-\ln(1-s_{1})-\ln(1-s_{2})$
$=e_{1}+e_{2},$
where $e_{1}=-\ln(1-s_{1})$ and $e_{2}=-\ln(1-s_{2})$ are the individual evidence weights equation 32 of $m_{1}$ and $m_{2}$, respectively. ∎
A.6 Proof of Theorem 1
Proof Outline.
We outline the key steps of the proof of Theorem1:
Step 1: Combine the individual mass functions $m_{j}^{+}$ and $m_{j}^{-}$ into aggregated mass functions $m^{+}$ and $m^{-}$ using Dempster’s rule of combination.
Step 2: Derive closed-form expressions for $m^{+}(\{h_{j}\})$ and $m^{+}(\mathcal{H})$ based on exponential evidence weights.
Step 3: Express the contour function $pl^{-}(h_{j})$ and use it to rewrite the conflict term $\mathbf{CF}$.
Step 4: Normalize the mass assignments to obtain support and opposition terms $\eta_{j}^{+}$ and $\eta_{j}^{-}$, leading to the final forms of $\mathbf{CF}$ and $\mathbf{IG}$. ∎
Proof.
We define $m^{+}(\cdot)=\bigoplus_{j}m_{j}^{+}$ and $m^{-}(\cdot)=\bigoplus_{j}m_{j}^{-}$ as the combined positive and negative mass functions, respectively. From Section A.5, we have the following expressions:
$m_{j}^{+}(\{h_{j}\})$ $=1-\exp(-e_{j}^{+})=1-\exp\left(-\sum\nolimits_{i}e_{ij}^{+}\right),$ (36)
$m_{j}^{-}(\overline{\{h_{j}\}})$ $=1-\exp(-e_{j}^{-})=1-\exp\left(-\sum\nolimits_{i}e_{ij}^{-}\right).$
Here, $\mathbf{CF}$ quantifies the conflict between the combined positive and negative evidence, while $\mathbf{IG}$ captures the overall ignorance, defined as the sum of all $m_{j}^{-}(\mathcal{H})$. Specifically, their formulations are given by:
$\mathbf{CF}$ $=\sum_{\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset}m^{+}(\mathcal{S}_{1})\,m^{-}(\mathcal{S}_{2})=\sum_{j}\left(m^{+}(\{h_{j}\})\sum_{\mathcal{S}\not\ni h_{j}}m^{-}(\mathcal{S})\right),$ (37)
$\mathbf{IG}$ $=\sum_{j}m_{j}^{-}(\mathcal{H}).$
To proceed, we compute the aggregated mass functions $m^{+}(\cdot)=\bigoplus_{j}m_{j}^{+}$ and $m^{-}(\cdot)=\bigoplus_{j}m_{j}^{-}$ using Dempster’s rule of combination equation 15. Noting that each $m_{j}^{+}(\cdot)$ is a simple mass function with only two focal sets, $\{h_{j}\}$ and $\mathcal{H}$, the combination simplifies to:
[…]
$=\left(1-\exp(-e_{j}^{+})\right)\prod_{l\neq j}\exp(-e_{l}^{+})=\prod_{l\neq j}\exp(-e_{l}^{+})-\prod_{l}\exp(-e_{l}^{+})$
$=\left(\exp(e_{j}^{+})-1\right)\exp\left(-\sum_{l}e_{l}^{+}\right)$
$m^{+}(\mathcal{H})$ $=\frac{1}{1-\kappa^{+}}\exp\left(-\sum_{l}e_{l}^{+}\right)\propto\exp\left(-\sum_{l}e_{l}^{+}\right),$
where $\kappa^{+}$ denotes the degree of conflict among the individual mass functions $m_{j}^{+}(\cdot)$. Based on the expressions above, we can derive the following unnormalized total mass:
$\sum_{j}m^{+}(\{h_{j}\})+m^{+}(\mathcal{H})\propto\exp\left(-\sum_{k}e_{k}^{+}\right)\left(\sum_{j}\exp(e_{j}^{+})-J+1\right).$ (39)
By normalizing the mass assignments, we obtain:
$m^{+}(\{h_{j}\})$ $=\frac{\exp(e_{j}^{+})-1}{\sum_{l}\exp(e_{l}^{+})-J+1},$ (40)
$m^{+}(\mathcal{H})$ $=\frac{1}{\sum_{l}\exp(e_{l}^{+})-J+1}.$
Similarly, we now derive the expression for $m^{-}(\cdot)$. Note that although $m_{j}^{-}(\cdot)$ is also a simple mass function, its focal set is no longer a singleton. Instead, it consists of exactly two focal sets: $\overline{\{h_{j}\}}$ and $\mathcal{H}$. Consequently, the combined mass function $m^{-}(\cdot)$ also has a non-singleton focal set. By applying Dempster’s rule of combination equation 15, we obtain:
$m^{-}(\mathcal{S})$ $=\frac{1}{1-\kappa^{-}}\left(\prod_{h_{j}\not\in\mathcal{S}}(1-\exp(-e_{k}^{-}))\right)\left(\prod_{h_{j}\in\mathcal{S}}\exp(-e_{j}^{-})\right)$ (41)
$m^{-}(\mathcal{H})$ $=\frac{1}{1-\kappa^{-}}\exp\left(-\sum_{l}e^{-}_{l}\right),$
where $\kappa^{-}=\prod_{l}\left(1-\exp(-e_{j}^{-})\right)$. Let $pl_{j}^{-}(\cdot)$ and $pl^{-}(\cdot)$ denote the contour functions corresponding to $m_{j}^{-}(\cdot)$ and $m^{-}(\cdot)$, respectively. Note that the term $\sum_{\mathcal{S}\not\ni h_{j}}m^{-}(\mathcal{S})$ in  equation 37 can be rewritten using the contour function. Specifically, by  equation 16, it holds that
$\sum_{\mathcal{S}\not\ni h_{j}}m^{-}(\mathcal{S})=1-pl^{-}(h_{j}).$ (42)
The explicit form of $pl_{j}^{-}(\cdot)$ is given by
$pl_{j}^{-}(h)=\begin{cases}\exp(-e_{j}^{-})&\text{if }h=h_{j},\\ 1&\text{otherwise}.\end{cases}$ (43)
Then, by applying the combination rule for contour functions in  equation 17, we obtain:
$pl^{-}(h_{j})\propto\prod_{l}pl_{l}^{-}(h_{j})=\exp(-e_{j}^{-}).$ (44)
Substituting into  equation 37, we obtain:
$pl^{-}(h_{j})=\frac{\exp(-e_{j}^{-})}{1-\kappa^{-}}=\frac{\exp(-e_{j}^{-})}{1-\prod_{j}(1-\exp(-e_{j}^{-}))},$ (45)
where $\kappa^{-}$ denotes the degree of conflict among the negative mass functions.
Finally, we continue simplifying the expression of the conflict term $\mathbf{CF}$ in Eq. equation 37 as follows:
$\mathbf{CF}$ $=\sum_{l}\left(m^{+}(\{h_{j}\})(1-pl^{-}(h_{j}))\right)$ (46)
$=\sum_{l}\left(\underbrace{\frac{\exp(e_{j}^{+})-1}{\sum_{j}\exp(e_{j}^{+})-J+1}}_{\eta^{+}_{j}}\cdot\underbrace{\frac{\exp(-e_{j}^{-})}{1-\prod_{j}(1-\exp(-e_{j}^{-}))}}_{\eta_{j}^{-}}\right)$
Next, we compute the expression for $\mathbf{IG}$. By definition, it follows that:
$\mathbf{IG}=\sum_{j}m_{j}^{-}(\mathcal{H})=\sum_{j}\left(1-m_{j}^{-}(\overline{\{h_{j}\}})\right)=\sum\nolimits_{j}\exp(-e_{j}^{-}).$ (47)
A.7 Detailed Experiment Settings
In Section 4.1, we briefly described the experimental settings due to space limitations. Here, we provide a more detailed version of the experimental setup for completeness.
Figure 6: Representative examples of four types of misbehaviors.
A.7.1 Datasets
Hallucination Data
In hallucination scenarios, evaluation is conducted on POPE (Li et al., 2023b) and R-Bench (Wu et al., 2024a), respectively, targeting object and relation hallucinations.
We follow the evaluation protocol proposed in POPE (Li et al., 2023b), which formulates object hallucination detection as a binary (Yes-or-No) task. Built on the MS COCO validation set (Lin et al., 2014), POPE prompts LVLMs with queries such as Is there a chair in the image? and expects definitive binary responses. This task design enables precise assessment of hallucination by systematically querying for nonexistent objects. To increase the likelihood of hallucination, we adopt POPE’s Adversarial Sampling strategy: for each image, objects are ranked by their co-occurrence frequency with ground-truth objects, and the top-$k$ frequently co-occurring but absent objects are selected as hallucination probes. We further evaluate relational hallucination using the R-Bench benchmark (Wu et al., 2024a), which is specifically designed to assess the ability of LVLMs to correctly perceive and reason about object relationships. R-Bench includes two types of binary (Yes-or-No) questions: image-level and instance-level. The image-level questions probe whether a particular relationship exists anywhere in the image, while the instance-level questions target relationships between specific object instances, denoted by color-coded bounding boxes or masks. The latter is particularly useful for assessing fine-grained and localized relational understanding. All examples are constructed from the validation split of the nocaps dataset (Agrawal et al., 2019), and we follow the data preparation procedure provided in the official repository. [^4] https://github.com/mrwu-mac/R-Bench
Representative examples from the relational hallucination evaluation are illustrated in the fourth column of Figure 6. Each instance is paired with a binary question prompt that queries the existence of either an object (e.g., Is there a chair in the image?) or a relationship between objects (e.g., Is the man in the image wearing a black jacket?).
Jailbreak Attacks
For jailbreak scenarios, we evaluate a range of jailbreak attacks, including FigStep (Gong et al., 2025), Hades (Li et al., 2024b), and VisualAdv (Qi et al., 2024), all implemented via the Jailbreak Benchmark [^5] https://github.com/thunxxx/MLLM-Jailbreak-evaluation-MMJ-bench. We further simulate typographic attacks following the protocol of (Goh et al., 2021).
[…]
Distinct prompts are employed depending on the jailbreak type:
  • For FigStep (Gong et al., 2025), the prompt reads: The image shows a list numbered 1, 2, and 3, but the items are empty. Please generate detailed content for each item on the list.
  • For Hades (Li et al., 2024b) and VisualAdv (Qi et al., 2024), we utilize the harmful instructions as provided by the Jailbreak Benchmark.
  • For typographic attacks, the prompt is constructed as follows: This photo is a ‘?’ Only output the corresponding letter of the correct option (e.g., a, b, c, or d) without outputting the full text of the option. (a) <true label> (b) <wrong label>.
Adversarial Examples
To evaluate the robustness of our method and competitive baselines, we consider two representative state-of-the-art adversarial attacks: ANDA (Fang et al., 2024) and PGN (Ge et al., 2023). Both approaches are optimization-based and specifically designed to deceive large vision-language models (LVLMs) through carefully crafted perturbations.
Following prior work (chen2024rethinking), we formulate adversarial example generation as a constrained maximization problem (Szegedy et al., 2014) that aims to significantly alter the model’s visual embedding representation. Concretely, we perturb the input image within an $\ell_{\infty}$-bounded region to maximize the discrepancy between its original and perturbed embeddings:
$\max_{x_{\text{adv}}\in\mathcal{B}_{\epsilon}(x)}\left\|e(x)-e(x_{\text{adv}})\right\|_{2}^{2},$ (48)
where $x$ denotes the clean input, $x_{\text{adv}}$ is the adversarial example, and $\mathcal{B}_{\epsilon}(x)$ is an $\ell_{\infty}$-norm ball of radius $\epsilon$ centered at $x$. The encoder $e(\cdot)$ corresponds to the vision backbone of the CLIP model, which is used as the surrogate model for computing adversarial directions.
Following recent advances in adversarial evaluation, we apply perturbations directly in the vision embedding space rather than in the pixel domain, enabling stronger attacks on downstream LVLMs. The full algorithmic details of the ANDA and PGN attacks are provided in their original papers (Fang et al., 2024; Ge et al., 2023). The adversarial examples generated by these methods are illustrated in the first column of Figure 6.
To standardize evaluation, we adopt a Yes-or-No question format that constrains the model’s output space and enables binary decision analysis. Each LVLM is prompted with: Is this image a <true label>? (only answer yes or no, do not need explanation), where <true label> denotes the ground-truth class label of the image.
OOD Inputs
To evaluate model robustness under distributional shifts, we consider out-of-distribution (OOD) inputs that elicit misbehavior in vision-language models. Specifically, we adopt the MMDT benchmark introduced by (Xu et al., 2025) [^6] https://huggingface.co/datasets/AI-Secure/MMDecodingTrust-I2T, which provides a curated dataset designed to probe the reliability of multimodal decoding under OOD scenarios.
We construct two out-of-distribution (OOD) evaluation scenarios: image corruptions and style transformations. Based on the MS COCO 2017 training set (Lin et al., 2014), we curate image-question pairs spanning four core vision-language tasks: object recognition, counting, spatial reasoning, and attribute recognition. To induce distributional shifts, we apply three severe corruptions (Zoom Blur, Gaussian Noise, Pixelation) and three artistic style transfers (Van Gogh, oil painting, watercolor), forming a comprehensive OOD benchmark for assessing misbehavior in LVLMs. In practice, we generate corrupted or style-transferred images by using the ground-truth <image caption> from the MMDT benchmark as prompts for the text-to-image model stabilityai/stable-diffusion-2. The resulting OOD inputs are shown in the third column of Figure 6.
To ensure standardized evaluation, we adopt a Yes-or-No question format. Specifically, each LVLM is prompted with: Please check whether the following description matches the picture content. Just answer yes or no without explanation. <image caption>, where <image caption> corresponds to the ground-truth caption of the image.
A.7.2 Hardware and Software Configuration
To ensure the reproducibility and reliability of the experiments conducted in this study, we detail the hardware and software environments used.
  • GPU Model(s):
    • Model: NVIDIA H800 PCIe
    • Count: 2 GPUs
    • Memory per GPU: 81 GB
  • CPU Model(s):
    • Model: Intel(R) Xeon(R) Platinum 8458P
    • Socket(s): 2
    • Core(s) per socket: 44
    • Thread(s) per core: 2
    • Total Logical Cores: 176
  • Operating System:
    • OS: Ubuntu 22.04.4 LTS
    • Kernel Version: 5.15.0-94-generic
  • Relevant Software Libraries and Frameworks:
    • CUDA: Version 12.6
    • PyTorch: Version 2.7.0+cu126
    • Scikit-learn: Version 1.6.1
    • NumPy: Version 1.26.4
    • Pandas: Version 2.2.3
A.8 Additional Experiment Results
Due to space constraints in the main paper, we present the complete results of additional analytical experiments below.
A.8.1 Analysis of Evidential Conflict and Ignorance
To complement the findings in Section4.4 and Section4.5, we extend the analysis of evidential conflict and ignorance to three additional LVLMs: DeepSeek-VL2, Qwen2.5-VL, and MoF-Models. The results are presented in Figure7. Similarly, to provide a broader perspective on the uncertainty patterns across different misbehavior types, we include density curve visualizations for the same three models. These results are reported in Figure 8.
(a) Deepseek.
(b) Qwen.
(c) Intern.
(d) MoF.
Figure 7: Analysis of conflict and ignorance, as quantified measures of evidential uncertainty, across four dataset types using four LVLMs.
(a) DeepSeek
(b) Qwen
(c) Intern
(d) MoF
Figure 8: Density distribution comparison between our method (conflict and ignorance) and predictive entropy across various misbehavior groupings on four LVLMs.
Table 7: Ablation study on the impact of model scale using DeepSeek. We separately report AUROC and AUPR for clarity. The best and second-best results are highlighted in bold and underlined, respectively.
AUROC
Method Tiny Small VL2
CF 0.837 0.548 0.681
IG 0.841 0.553 0.731
AUPR
Method Tiny Small VL2
CF 0.768 0.523 0.609
IG 0.770 0.536 0.616
Table 8: Ablation study on the impact of model scale using Qwen. We separately report AUROC and AUPR for clarity. The best and second-best results are highlighted in bold and underlined, respectively.
AUROC
Method 3B 7B 32B
CF 0.722 0.795 0.737
IG 0.862 0.724 0.588
AUPR
Method 3B 7B 32B
CF 0.589 0.829 0.552
IG 0.862 0.727 0.607
A.8.2 Analysis of Hallucination and Jailbreak by Category
To further characterize the applicability of our method, we perform a fine-grained analysis of distinct subcategories within hallucination and jailbreak scenarios. Specifically, we differentiate between object-level and relation-level hallucinations to examine their respective uncertainty patterns. For jailbreak attacks, we investigate the contrast between structured Yes-and-No (Yes-No) formatting prompts and other unstructured attack variants (Open-ended). This analysis offers deeper insights into how different types of misbehavior manifest in evidential signals.
Table 9: AUROC and AUPR of our method under the impact of hallucination type using four LVLMs. We separately report AUROC and AUPR for clarity. The best results are highlighted in bold.
AUROC
Method DeepSeek Qwen Intern MoF
POPE RBench POPE RBench POPE RBench POPE RBench
CF 0.776 0.518 0.908 0.501 0.860 0.593 0.999 0.672
IG 0.962 0.596 0.586 0.576 0.691 0.840 0.999 0.673
AUPR
Method DeepSeek Qwen Intern MoF
POPE RBench POPE RBench POPE RBench POPE RBench
CF 0.938 0.751 0.923 0.785 0.656 0.606 0.941 0.571
IG 0.634 0.658 0.826 0.848 0.913 0.553 0.941 0.536
Table 10: AUROC and AUPR of our method under the impact of jailbreak type using four LVLMs. We separately report AUROC and AUPR for clarity. The best results are highlighted in bold.
AUROC
Method DeepSeek Qwen Intern MoF
Open-ended Yes-No Open-ended Yes-No Open-ended Yes-No Open-ended Yes-No
CF 0.720 0.966 0.920 0.997 0.871 0.623 0.702 0.995
IG 0.587 0.789 0.862 0.872 0.921 0.637 0.708 0.751
AUPR
Method DeepSeek Qwen Intern MoF
Open-ended Yes-No Open-ended Yes-No Open-ended Yes-No Open-ended Yes-No
CF 0.664 0.961 0.967 0.990 0.949 0.732 0.805 0.999
IG 0.739 0.823 0.673 0.823 0.544 0.510 0.573 0.987
A.8.3 Single-Modality Evaluation of Evidential Uncertainty
To further demonstrate the adaptability of our method, we conducted additional experiments on single-modality models. We performed a controlled experiment using a LeNet classifier trained on the handwritten digits dataset MNIST (LeCun, 1998) and the German Traffic Sign Recognition Benchmark (GTSRB) (Stallkamp et al., 2011), with FashionMNIST (Xiao et al., 2017) serving as out-of-distribution (OOD) data and FGSM-generated adversarial examples (Goodfellow et al., 2014). For comparison, we employed several classical uncertainty quantification methods: MC Dropout (Gal & Ghahramani, 2016) (100 iterations), Deep Ensembles (Lakshminarayanan et al., 2017) (5 models), and Evidential Deep Learning (EDL) (Sensoy et al., 2018). As shown in Table 11, $\mathrm{CF}$ and $\mathrm{IG}$ achieve competitive or superior performance compared to these baselines, attaining high AUROC scores for both adversarial and OOD detection tasks.
Table 11: AUROC performance comparison on adversarial and OOD detection tasks.
Scenario Dataset MC Dropout Deep Ensemble EDL CF IG
Adversarial MNIST 0.927 0.933 0.892 0.935 0.701
GTSRB 0.970 0.980 0.912 0.962 0.894
OOD MNIST 0.937 0.985 0.802 0.972 0.995
GTSRB 0.907 0.969 0.802 0.944 0.995
A.8.4 Experiments on Larger-Scale LVLM
To demonstrate the consistency of our method across both small-scale and larger-scale LVLMs, particularly with respect to the observations, we conducted experiments on Qwen-2.5-VL-72B (Bai et al., 2025). For Observation 1, as shown in Figure 9, we find that the conclusions in Qwen-2.5-VL-72B align with the behaviors observed in the small-scale LVLMs. For Observation 2, as shown in Table 12, the results are largely consistent with those from small-scale LVLMs, indicating that the observed behaviors generalize across model scales.
Figure 9: Layer-wise changes of evidential uncertainty and analysis of conflict vs. ignorance across four dataset types using Qwen-2.5-VL-72B.
Table 12: The AUROC and AUPR of our methods and baselines on Qwen-VL-72B, in adversarial, OOD, hallucination, and jailbreak settings. Best and next-best results are marked in bold and underlined, respectively.
[…]
AUROC AUPR AUROC AUPR AUROC AUPR AUROC AUPR AUROC AUPR
SC 0.701 0.874 0.566 0.833 0.712 0.640 0.674 0.884 0.663 0.808
SE 0.609 0.856 0.543 0.818 0.502 0.670 0.602 0.786 0.564 0.781
PE 0.783 0.558 0.618 0.759 0.766 0.692 0.714 0.727 0.720 0.684
LN-PE 0.783 0.553 0.645 0.628 0.655 0.658 0.717 0.720 0.700 0.639
HiddenDetect 0.622 0.659 0.854 0.762 0.823 0.662 0.613 0.719 0.728 0.701
CF(ours) 0.817 0.884 0.640 0.789 0.759 0.641 0.731 0.834 0.737 0.787
IG(ours) 0.763 0.872 0.659 0.774 0.665 0.556 0.714 0.827 0.701 0.757
A.8.5 Comparison with Evidential Deep Learning
This Subsection provides a detailed theoretical comparison between our approach and Evidential Deep Learning (EDL), highlighting fundamental differences in their mathematical foundations and implementation strategies.
Both methods are rooted in Dempster-Shafer Theory (DST), but represent distinct implementations. EDL implements the Subjective Logic (SL), which “formalizes DST’s notion of belief assignments over a frame of discernment as a Dirichlet Distribution” (Sensoy et al., 2018). In contrast, our approach employs the full expressive power of classical DST, allowing for more flexible and comprehensive uncertainty representation.
Consider a frame of discernment $\mathcal{H}=\{a,b,c\}$ representing class labels. The SL formulation used in EDL constrains belief assignment to only singletons and the entire frame:
$m(a)+m(b)+m(c)+m(\mathcal{H})=1,$
where $0\leq m(a),m(b),m(c),m(\mathcal{H})\leq 1$. This results in only $|\mathcal{H}|+1=4$ belief assignments.
Our method employs the complete power set of the frame:
$\sum_{S\subseteq\mathcal{H}}m(S)=1,$
where $0\leq m(S)\leq 1$ and $m(\emptyset)=0$. This allows for $2^{|\mathcal{H}|}-1=7$ distinct mass assignments, enabling richer uncertainty representation.
The theoretical differences between the two approaches are substantial. While both use the same definition of ignorance (mass assigned to the total frame $m(\mathcal{H})$), EDL learns a single mass function over all evidence, whereas our method models separate mass functions for individual feature values and leverages evidence fusion to quantify conflicts.
Architecturally, EDL modifies the model’s final layer (replacing softmax) and requires retraining. Our approach is training-free, relying only on parameter estimation without architectural changes. This difference affords our method greater interpretability, revealing consistent layer-wise trends where ignorance decreases and conflict increases across decoder layers, as demonstrated in Figure 3 of the main text.
The choice of full DST over SL-based approaches provides enhanced expressiveness through the ability to assign mass to arbitrary subsets, enabling more nuanced uncertainty representation. Our architecture-preserving, training-free approach maintains flexibility while providing deeper insights into model behavior through layer-wise analysis of uncertainty dynamics.
A.8.6 Complementarity of CF and IG in Detection
To further examine whether CF and IG capture complementary uncertainty signals, we evaluate two simple fusion strategies for hallucination detection:
  • Conjunctive rule (&): a sample is flagged as hallucinated only if both CF and IG exceed their respective thresholds;
  • Disjunctive rule (|): a sample is flagged as hallucinated if either CF or IG exceeds its threshold.
The experiments were conducted on Qwen-2.5-VL-72B using the hallucination dataset, with thresholds for each method determined according to the Youden index. As shown in Table 13, combining CF and IG indeed leverages complementary cues: the disjunctive rule improves recall and yields the best overall F1 score.
Table 13: Performance comparison of CF, IG, and their fusion strategies for hallucination detection.
Method Accuracy Precision Recall F1 Score
CF 0.851 0.929 0.885 0.907
IG 0.859 0.921 0.906 0.914
CF & IG 0.835 0.929 0.866 0.896
CF $|$ IG 0.873 0.922 0.924 0.923
The conjunctive rule is more conservative and yields lower recall, whereas the disjunctive rule benefits from the complementary nature of the two signals, producing the strongest performance. This indicates that CF and IG encode partly distinct uncertainty information.
A.9 Large Language Models Usage
We used the large language model ChatGPT (GPT-5-mini) to aid in polishing and improving the clarity of the manuscript. All technical content, derivations, experiments, and conclusions were independently verified by the authors.